# Modular Compilation for Quantum Chiplet Architectures

### v3

@@ Abstract.
As quantum computing technology matures, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets. However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessarily support a universal basis gate set. Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization, are typically not designed for qubit links of significantly varying latency or fidelity. In this work, we propose SEQC, a hierarchical parallelized compilation pipeline optimized for chiplet-based quantum systems, including several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC attains a $9.3\%$ average increase in circuit fidelity (up to $49.99\%$). Additionally, owing to its ability to parallelize compilation, SEQC achieves 3.27×3.27$\times$3.27 × faster compilation on average (up to 6.74×6.74$\times$6.74 ×) over a chiplet-unaware Qiskit baseline.
@@ 1. Introduction
We are already observing signs of this shift. Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling (smith_2022_chiplets). Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips (Field_2024_RigettiModular). Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2023_Roadmap; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
@@
Unfortunately, current compilation infrastructure does not support quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution [^1] The term “execution” here includes all shots required to produce a final result for the user. Different executions of the same circuit can span qubit calibration windows., as the ideal physical qubits to target change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in a compilation complexity that is quadratic to the number of qubits in the quantum processor (experimentally observed in Figure 12).
Hardware modularity adds significant complexity to this already hard task, as inter-chiplet links are typically inferior compared to intra-chiplet ones (laracuente2022modeling; smith_2022_chiplets), connectivity across chiplets is often limited (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers), and not all basis gates are necessarily supported across chiplets (niu2023low). In fact, popular quantum software stacks today (e.g., Qiskit (qiskit2024)) are not even cognizant of the existence of hardware modularity. Hardware modularity, though, also presents an opportunity. Inspired by classical compilation, in this paper we leverage hardware modularity to achieve compilation modularity.
@@
Compilation in classical systems is typically performed independently and in parallel for each source file, producing one object file per source. The individual object files are then linked together to construct the executable. We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each chiplet, and maps subcircuits to chiplets, and then in parallel elaborates each subcircuit and compiles it for its target chiplet. This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given chiplet architecture, and performs only the elaboration step recurrently before each execution.
@@
As the qubit capacity $n$ of quantum processors grows exponentially, the $O(n^{2})$ compilation latency of contemporary compilers rises even faster. SEQC replaces the recurrent $O(n^{2})$ compilation step with several parallel $O(k^{2})$ elaboration steps for $k$-qubit chiplets. The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once. The recurrent compilation in SEQC is only $O(k^{2})$, and as the number of qubits $k$ per chiplet is expected to remain relatively stable, or grow at a much slower pace in the foreseeable future (IBM_2023_Roadmap), SEQC’s recurrent compilation latency will barely grow as processors scale.
@@
Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to chiplets to minimize inter-chiplet communication. As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-chiplet gates compared to today’s stock compilers, leading to higher execution fidelity. More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
@@
  • We make stock compilers aware of hardware modularity, thereby allowing them to correctly compile circuits for modular architectures with limited cross-chiplet gate support.
  • We design and implement SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or chiplet splitting) performed only once for a given architecture, and the second stage (elaboration, or chiplet compilation) performed in parallel for each chiplet. Only this second stage needs to be performed before each execution, and thus SEQC’s compilation time is largely unaffected by the growth of qubit counts in future quantum processors.
  • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC compiles circuits with $9.3\%$ higher circuit fidelity on average (up to $49.99\%$), while consistently achieving 3.27×3.27$\times$3.27 × faster compilation time on average (up to 6.74×6.74$\times$6.74 ×) compared to a chiplet-unaware Qiskit baseline.
@@ 2. Background and Motivation
(a) Example 2-chiplet architecture
(b) Input circuit from qubit allocation
(c) Insertion of perfect-shuffle

### v4

# ab,ab.braket

@@ Abstract.
[…]
@@ 1. Introduction
We are already observing signs of this shift. Recent developments in quantum chip linking, including flip-chip architectures (gold2021entanglement) and low-loss coaxial cables (niu2023low), suggest that modular designs are the most viable for scaling (smith_2022_chiplets). Many contemporary or upcoming leading quantum systems adopt chiplet-based modular quantum processors, for example using high-bandwidth quantum links between nearest-neighbor quantum chiplets (e.g., carrier-chip couplers in Rigetti Aspen-M (Rigetti_AspenM; gold2021entanglement), or m-couplers in IBM Crossbill (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or lower-fidelity, lower-bandwidth, but longer-distance flexible coupling of discrete chips (e.g., l-couplers in IBM Flamingo (gambetta_2024_IBMcouplers; Mandelbaum_2024_IBMCouplers)), or a combination of the above (e.g., c-, m-, and l-couplers in IBM Starling (Mandelbaum_2024_IBMCouplers)), or multi-chip connectivity through tunable couplers and routing chips in Rigetti Cepheus (Rigetti_Cepheus; Field_2024_RigettiModular). Most major companies have set their sights at modular designs to meet quantum scaling targets in practical ways, and modular quantum processors dominate their latest roadmaps (IBM_2025_FTQC; IonQ_2024_Roadmap; Rigetti_2024_Roadmap).
At the same time, no single qubit technology simultaneously offers long coherence, fast gates, high connectivity, and scalable, mature fabrication. Superconducting qubits offer very fast gate latencies and mature fabrication, but low coherence times and mostly local connectivity. Trapped ions have excellent coherence but slow speeds, and all-to-all connectivity but only for the few ions located at the C-cut sapphire window. Neutral atoms have excellent scalability and arrangeable connectivity, but do not reach the gate speeds and manufacturing robustness of superconducting qubits, or the coherence times of trapped ions. Hybrid, modular quantum architectures, comprising QPUs based on different qubit modalities, provide a path to combine the complementary strengths of each platform, enabling performance impossible for any single modality. For example, superconducting qubits can serve as fast computational QPUs, while ions can provide memory QPUs, and Rydberg atoms can provide ancillae and distillation functionality. Such heterogeneity mirrors the classical transition from monolithic CPUs to systems composed of CPUs, GPUs, NPUs, and specialized accelerators. As such, hybrid modular quantum systems are increasingly recognized as a pragmatic design path for long-term universal quantum computers (DARPA_HARQ; HQAN; hetarch; Xiang_superconducting_hybrid_2013). Efforts to realize hybrid quantum systems have predominantly leveraged quantum state transfer, i.e., SWAPs, between heterogeneous qubit technologies (Kurizki_hybrid_2015; Clerk2020; Zhou2014; PhysRevLett.108.130504). While universal quantum operations, e.g., CNOT or CZ operations, that interface two different quantum technologies have been demonstrated (PhysRevLett.114.080501; Xiang_superconducting_hybrid_2013; Cia_2024; Daniilidis_2013; Yu2016), these often have to compromise on gate fidelity, and have yet to be realized for all combinations of common qubit technologies, e.g, between superconducting and trapped-ion qubits. Thus, hybrid quantum systems may exhibit interconnect heterogeneity, where inter-module links may not be able to support universal gate sets, while intra-module couplers do.
@@
Unfortunately, current compilation infrastructure does not support quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution [^1] The term “execution” here includes all shots required to produce a final result for the user. Different executions of the same circuit can span qubit calibration windows., as the ideal physical qubits to target change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in a compilation complexity that is quadratic to the number of qubits in the quantum processor (experimentally observed in Figure 11(a)). Hardware modularity, though, also presents an opportunity. Inspired by classical compilation, in this paper we leverage hardware modularity to achieve compilation modularity.
@@
Compilation in classical systems is typically performed independently and in parallel for each source file, producing one object file per source. The individual object files are then linked together to construct the executable. We propose a compilation framework for modular quantum processors that works in a similar fashion: it stratifies, i.e., splits, the source quantum circuit into subcircuits small enough to fit in each module, and maps subcircuits to modules, and then in parallel elaborates each subcircuit and compiles it for its target module. This Stratify-Elaborate Quantum Compiler (SEQC) stratifies a source program only once for a given module architecture, and performs only the elaboration step recurrently before each execution.
@@
As the qubit capacity $n$ of quantum processors grows exponentially, the $O(n^{2})$ compilation latency of contemporary compilers rises even faster. SEQC replaces the recurrent $Oa(n^{2})$ compilation step with several parallel $O(k^{2})$ elaboration steps for $k$-qubit modules. The stratification step is $O(n^{2})$, so the end-to-end complexity remains the same, but stratification is performed only once. The recurrent compilation in SEQC is only $O(k^{2})$, and as the number of qubits $k$ per module is expected to remain relatively stable, or grow at a much slower pace in the foreseeable future (IBM_2025_FTQC), SEQC’s recurrent compilation latency will barely grow as processors scale.
@@
Additionally, as SEQC is cognizant of hardware modularity, it can stratify the source quantum circuit into subcircuits and map them to modules while minimizing cross-module communication. As we show in this paper, SEQC produces circuits with shorter execution times and significantly fewer inter-module gates compared to today’s stock compilers, leading to higher execution fidelity. More importantly, as the number of qubits in a processor grows, SEQC achieves even higher performance in these figures of merit.
@@
  • We make stock compilers aware of modularity through a “peephole” correction pass, thereby allowing them to correctly compile circuits for modular architectures with limited cross-module gate support.
  • We design and implement SEQC, a Stratify-Elaborate Quantum Compiler for modular architectures. SEQC performs compilation in two stages, with the first stage (stratification, or module splitting) performed only once for a given architecture, and the second stage (elaboration, or module compilation) performed in parallel for each module. Only this second stage needs to be performed before each execution, and thus SEQC’s compilation time is largely unaffected by the growth of qubit counts in future quantum processors.
  • We design and implement in SEQC several novel methods for qubit placement, qubit routing, and circuit optimization. In particular, we classify swaps into symbiotic, commensalistic, and parasitic, based on how swap placement affects inter-module gates, and develop algorithms to promote symbiotic and avoid parasitic swap placement; we develop qubit “pinning” techniques to allow elaboration without disrupting module connectivity; and we adapt Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) techniques to aid in circuit stratification.
  • We evaluate SEQC on a modular quantum system consisting of small-scale (10-qubit heavy-hex) modules, and find that it compiles circuits with $9.3\%$ higher circuit fidelity on average (up to $49.99\%$), while consistently achieving faster compilation times ($3.27\times$ on average, up to $6.74\times$) compared to a module-unaware Qiskit baseline. When evaluated on a system with large-scale (120-qubit grid topology) modules, SEQC achieves $23.3\%$ higher fidelity on average ($63.36\%$ max), while compiling $1.34\times$ faster on average (up to $3.37\times$). More importantly, SEQC outperforms Qiskit even when the latter is afforded significant allowances (no universal gate restrictions, and much lower inter-module gate errors).
@@ 2. Background and Motivation
Compilation is often divided into distinct qubit allocation (layout and routing), basis translation, and optimization stages.
Qubit Allocation
Qubit allocation addresses the topology constraints of quantum devices, where physical qubits on a device have limited connectivity to other qubits on the device. It is critical that after this stage, which often requires the modification or addition of circuit gates, the resulting optimized, technology-dependent circuit is functionally equivalent to the original, technology-independent algorithm (smith2019quantum). As qubit allocation is NP-complete (qubit_allocation), it is typically divided into two substages, layout and routing, to make it more tractable.
Logical-to-physical qubit layout, also known as qubit placement, involves deriving an injective map from the virtual qubits of the quantum circuit to the physical qubits of the quantum device (qubit_routing; sabre_swap), cognizant of its limited connectivity. In other words, the qubit layout stage decides the initial position of logical qubits on the device.
Qubit routing operates from an initial qubit layout and inserts SWAP operations to align a quantum circuit’s gates to a given qubit topology. Consider, for example, the state-of-the-art in qubit allocation: the SABRE heuristic search algorithm (sabre_swap). In SABRE, the distinction of the layout and routing stages motivates an iterative method for solving qubit allocation: an initial layout is used to perform qubit routing, which, in turn, is used to generate a new initial layout. Because qubit routing remains NP-hard (qubit_routing), the problem is often simplified by assuming qubit links are roughly equivalent, particularly with respect to average fidelity and two-qubit gate duration. This assumption can manifest implicitly when a given device topology is represented as an unweighted, undirected graph. For monolithic architectures, one could debate the marginal tradeoff between performance and solution quality. However, for modular architectures, the cross-module connections are expected to be far worse compared to intra-module connections (smith_2022_chiplets; PhysRevLett.114.080501; Zhou2014; PhysRevLett.108.130504; Yu2016). Thus, as we will show, the current simplified model applied on modular architectures leads to worse solutions than a more physically accurate one.
(a) Example 2-module
architecture
(b) Input circuit from
qubit allocation
(c) Insertion of
perfect-shuffle SWAPs